OpenStudy (anonymous):

please help me with this problem : find the condition for the equations ax^2+bx+c=0 and a'x^2+b'x+c'=0 have reciprocal roots.

OpenStudy (raden):

equations ax^2+bx+c=0 and a'x^2+b'x+c'=0 have reciprocal roots, if the product all roots must be equal 1. prove : c/a * a'/c' = c/a * a/c = 1 QED (^_*)

OpenStudy (anonymous):

if p and q be the roots of ax^2+bx+c=0 then p+q = -b/a and p.q = c/a as the roots are reciprocal then the roots of a'x^2+b'x+c'=0 will be 1/p and 1/q ?? hence 1/p + 1q = =b'/a' or (p+q)/pq = b'/a' and 1/p.1/q = c'/a' or 1/p.q = c'/a' ?

OpenStudy (raden):

a'x^2+b'x+c' a' and c' here means you have to replace a and c, so the equation becomes cx^2 + bx + a = 0

OpenStudy (anonymous):

then how do i find the condition that they have reciprocal roots as asked in the question?

OpenStudy (raden):

like i said above, read my first coment

OpenStudy (anonymous):

the answer given is b/b' = c/a'= a/c', i am not getting you what is wrong in this process if p and q be the roots of ax^2+bx+c=0 then p+q = -b/a and p.q = c/a as the roots are reciprocal then the roots of a'x^2+b'x+c'=0 will be 1/p and 1/q ?? hence 1/p + 1q = =b'/a' or (p+q)/pq = b'/a' and 1/p.1/q = c'/a' or 1/p.q = c'/a' ?

OpenStudy (raden):

too long, but i see that. (btw, 1/p + 1/q = -b'/a')

OpenStudy (anonymous):

the answer given is b/b' = c/a'= a/c' how does it come ?

OpenStudy (raden):

let the roots of ax^2 + bx + c = 0 are p and q we know that the sum of roots and the product of roots above is : p+q = -b/a p * q = c/a respectively. and supposed the roots of a'x^2 + b'x + c' = 0 are p' and q' we know that the sum of roots and the product of roots above is : p'+q' = -b'/a' p' * q' = c'/a' now will be proved that b/b' = c/a'= a/c' known that p' = 1/p, and q' = 1/q p'q' = 1/p * 1/q = 1/pq c'/a' = 1/(c/a) = a/c or => cc' = aa' => c/a' = a/c' p'+q' = 1/p + 1/q -b'/a' = (p+q)/pq -b'/a' = (-b/a)/(c/a) = -b/c ............................................... x (-1) b'/a' = b/c or =>b'c = ba' => c/a' = b/b' therefore, b/b' = c/a'= a/c' QED

OpenStudy (anonymous):

fantastic .. thanx a lot for your step by step explanation :)

OpenStudy (raden):

YOU'RE WELCOME :)